19. Project Analysis (LO1, 2, 3, 4] You are considering a new product launch. The project will cost $1,700,000, have a four-y please help me.. i dont know what to do with this “unit sales vc fc are probably accurate with the +- 10%”.. please help me by giving the precise amount of best worst base case.. thanks
Answer #1

Given,

Project Cost = $ 1,700,000

Project life = 4 years

Salvage Value = 0

Depreciation = 0

Projected Sales = 190 units per year

Price per Unit = $18,000

Variable Cost per unit = $11,200

Fixed Cost = $ 410,000 Per Year

Tax = 35%

Required rate of Return (K) = 12%

a.

Base Case (BC) Upper Bound (UB) Lower Bound (LB)
Projected Sales 190.00 209.00 171.00
Price per Unit $18,000 $19,800.00 $16,200.00
Variable Cost per unit $11,200.00 $12,320.00 $10,080.00
Fixed Cost $410,000.00 $451,000.00 $369,000.00

Upper Bound (UB) = BC * 1.1 --> + 10%

Lower Bound (LB) = BC * 0.9 --> - 10%

Sales $3,420,000.00
Variable Cost $2,128,000.00
Fixed Cost $410,000.00
Depreciation 0
EBIT $882,000.00
Tax 35.00%
Net Income $573,300.00

Sales = Price per Unit * Project Sales

Variable cost = Varible cost per unit * Projected Sales

EBIT = Sales - Variable Cost - Fixed Cost - Depreciation

Net Income = EBIT - (EBIT*Tax)

Cash flow = EBIT + Depreciation - Taxes

= $882,000.00 + 0 - $308,700.00

= $573,300.00

NPV Base case = Investment + Cash Flow * ((1 - (1 + K)^-t)/k)

= $ 1,700,000 + $573,300.00 * ((1 - (1 + .12)^-4) / 0.12

= $ 1,700,000 + $573,300.00 * (0.36 / 0.12)

= $ 1,700,000 + $573,300.00 * 3.037

= $ 3,441,312.4

Worst Case Best Case
Projected Sales 171.00 209.00
Price per Unit $19,800.00 $16,200.00
Variable Cost per unit $12,320.00 $10,080.00
Fixed Cost $451,000.00 $369,000.00
Depreciation 0 $0.00
Total Sale Value $3,385,800.00 $3,385,800.00
EBIT $2,922,480.00 $3,006,720.00
Tax 35.00% 35.00%
Net Income $1,899,612.00 $1,954,368.00
Case Flow $1,899,612.00 $1,954,368.00
NPV $7,469,785.27 $7,636,098.37

C. Cash Break-Even = (FC - Depreciation) / P - V

= ($410,000.00 - 0) / ($18,000 - $11,200) = $410,000.00 / $ 6,800 = 60.29

d.

Accounting Break-Even = (FC) / P - V

= ($410,000.00) / ($18,000 - $11,200) = $410,000.00 / $ 6,800 = 60.29